OpenStudy (anonymous):

Eliminate the parameter to find a Cartesian equation of the curve. x= squrt t+1, y=squrt t-1

OpenStudy (anonymous):

dose any one know the answer ?

OpenStudy (campbell_st):

well are the equations \[x = \sqrt{t + 1}...and.....y = \sqrt{t -1}\]

OpenStudy (anonymous):

yep

OpenStudy (campbell_st):

ok... so looking at the x equation can't you say \[x^2 -1 = t\] then substitute that into the y equation...